# SimCity: Multi-Agent Urban Development Simulation with Rich Interactions

### v2

@@ Abstract
Large Language Models (LLMs) open new possibilities for constructing realistic and interpretable macroeconomic simulations. We present SimCity, a multi-agent framework that leverages LLMs to model an interpretable macroeconomic system with heterogeneous agents and rich interactions. Unlike classical equilibrium models that limit heterogeneity for tractability, or traditional agent-based models (ABMs) that rely on hand-crafted decision rules, SimCity enables flexible, adaptive behavior with transparent natural-language reasoning. Within SimCity, four core agent types (households, firms, a central bank, and a government) deliberate and participate in a frictional labor market, a heterogeneous goods market, and a financial market. Furthermore, a Vision–Language Model (VLM) determines the geographic placement of new firms and renders a mapped virtual city, allowing us to study both macroeconomic regularities and urban expansion dynamics within a unified environment. To evaluate the framework, we compile a checklist of canonical macroeconomic phenomena, including price elasticity of demand, Engel’s Law, Okun’s Law, the Phillips Curve, and the Beveridge Curve, and show that SimCity naturally reproduces these empirical patterns while remaining robust across simulation runs.
footnotetext: * Equal contribution.footnotetext: $\dagger$ Corresponding authors.
@@ 1 Introduction
For the past two decades, the Dynamic Stochastic General Equilibrium (DSGE) framework has been the predominant paradigm for studying aggregate economic behavior (Blanchard, 2009; Glandon et al., 2023). While mathematically elegant, DSGE relies on the assumption of simplified representative agents solving explicit optimization problems, which restricts heterogeneity and richer behavioral dynamics (Sergi, 2018; Vines and Wills, 2020; Storm, 2021). Agent-based models (ABMs) offer a bottom-up alternative that accommodates heterogeneity but typically rely on hand-crafted decision rules (Dilaver et al., 2018). Advancing the ABM tradition with recent developments in LLMs, we propose SimCity, a multi-agent macroeconomic simulation framework that enables flexible, adaptive behavior with transparent natural language reasoning, and a virtual city environment that provides spatial context.
SimCity models the economy as the interaction of four types of agents: households, firms, a central bank, and a government. To incorporate realistic heterogeneity, households and firms are instantiated as families of agents with rich variation in preferences, abilities, and other background, whereas the central bank and government are modeled as single institutional agents, in line with the economic literature (Blanchard, 2025). Agents are implemented through an LLM-based module that integrates environmental observations, agent-specific traits, and structured memory to generate reasoning, planning, and decision-making. Following the classic agent–interaction–environment paradigm of multi-agent system design (Wooldridge, 2009), we construct a simulation environment that features a frictional labor market, heterogeneous goods markets, and core financial interactions. To provide spatial context for both agents and researchers, the environment is augmented with a visualized map (Figure 1).
Figure 1: The framework of SimCity. Left: A visualized map with three types of buildings. Right: The rich interactions between various agent modules.
@@
1. LLM-driven macroeconomic simulation with rich interactions. We integrate LLMs into economic simulation by modeling four distinct economic roles, households, firms, a central bank, and a government, as LLM-driven agents. This enables realistic agent heterogeneity and rich economic interactions.
@@
3. Systematic evaluation. We compile a comprehensive checklist of canonical macroeconomic phenomena and regularities and use it to demonstrate that our framework exhibits established macroeconomic patterns and generates plausible responses to exogenous shocks.
@@ 2 Related Work
DSGE and ABM represent two leading, and in many respects, contrasting approaches to macroeconomic modeling. DSGE features highly idealized decision-makers characterized by precise mathematical objective functions, rational expectations about future variables, and strict market-clearing equilibrium conditions (Stokey et al., 1989; Ljungqvist and Sargent, 2018). In contrast, ABM follows a bottom-up, simulation-driven methodology, where agents are endowed with rule-of-thumb behavior patterns and, in some cases, learning heuristics (Tesfatsion and Judd, 2006; Axtell and Farmer, 2025). Although DSGE remains the dominant approach in economics departments and policy institutions (Woodford, 2009; Negro et al., 2013; Chen et al., 2023), there has been a renewal of interest in ABMs, particularly following the Global Financial Crisis of 2008 and again after the COVID-19 crisis, as mainstream DSGE models struggled to account for these unprecedented shocks (Gatti et al., 2008; Stiglitz, 2018; Borsos et al., 2024). Our work is closer to the ABM tradition. However, we depart from conventional ABMs by replacing predetermined behavioral rules with LLM-powered agents capable of flexible reasoning and information processing in natural languages.
Li et al. (2024) introduces EconAgent as a recent attempt at LLM-driven macroeconomic simulation. Its framework restricts agent-environment interactions to two simplified variables: consumption share and work propensity, which constrains the range of economic activities that can be simulated. Our work extends this approach by incorporating other key economic elements, including firms as LLM-driven agents, heterogeneous goods markets, enriched financial activities, and a taxation system. Thus, our framework allows the exploration of a wider range of macroeconomic phenomena. Another contemporary work Mi et al. (2025), which build similar cluster of agents, is proposed as a testbed for performance evaluation while we focus on urban development and macroeconomic laws.
Simulator AI Econ. LEN CATS EconAgent SimCity
@@
Tax Type Labor $\times$ $\times$ Labor Labor&
VAT
@@
Table 1: A comparison of agent-driven macroeconomic simulators.
Regularity AI- Econ. Econ Agent LEN CATS SimCity (Ours)
Phillips Curve / $\times$ $\times$
Okun’s Law /
Beveridge Curve / / $\times$ /
Price Elasticity of Demand / / / /
Engel’s Law / / / /
Investment Volatility / / / /
Price
Stickiness / / / /
Legend: ✓: Verified; $\times$: Inconsistent; /: Not available.
Note: The LEN paper claims verified the Beveridge Curve, but we are unable to reproduce it with https://github.com/newwayland/baseline-economy
Table 2: Checklist of macroeconomic regularities considered in this work.
@@ 3.1 Environment
In phase 2 (the development phase), the population is fixed. However, urban developments such as firm creation, investment, and spatial reallocation, continue to take place. We focus on this phase to study macroeconomic regularities, as the system operates in a quasi-steady state that is less susceptible to the transient noise present during the expansion period.
Each phase lasts some steps. Each simulation step represents one month. The order of events within each step is detailed in Appendix D.2. By design, agents can only access information from previous steps.
@@
We develop a web-based render module to visualize the urban expansion as shown in Figure 1. All buildings are displayed on the rendered map. The geographic placement of new firms is decided by a Visual-Language Model (VLM). For technical details of the module, please refer to Appendix D.3.
@@ 3.2 Interaction Protocol
We leverage the common-sense reasoning capabilities of LLMs to act as human-like, heterogeneous agents. Details about the prompts used are provided in Appendix E.1. Agents interact with the environment by means of function calling. The framework loads all the operations that the agent can execute and appends formatted function names along with their descriptions to the prompt. The LLMs will return the actions to be taken and their parameters in JSON format. The framework will execute after a verification. For detailed examples, see Appendix E.3.
@@ 4.1 Households
Each month, households receive two types of information: a personal report summarizing income, expenditures, and other status changes during the current period, and a citywide update containing goods prices, labor market conditions, housing availability, returns on investment, and interest rates. Additionally, households may receive offers for vacancies based on their skill alignment.
@@ 4.2 Investment Pool and Firms
Each firm in SimCity produces a single type of good. It hires households as workers, invests in capital to improve productivity, and transforms input goods to its specialized output, which is then sold in the market. Firms are instantiated from templates that specify skill requirements of job positions in the firm and the input-output relations governing its production process, where a firm may take possibly more than one input good for its production. The effectiveness of resource utilization (i.e., the ratio at which input goods are transformed into the output good) is determined by a firm-specific Cobb-Douglas production function (Cobb and Douglas, 1928):
@@
where $A$ is total factor productivity, $K$ is the firm’s capital stock, $L$ is the effective labor supply, defined as the amount of labor input adjusted for the match between job skill requirements and employees’ actual skills.
@@ 4.3 Government
$t_{i}=\sum_{k=1}^{K}r_{k}\,\bigl[\,\min\{\,z_{i},\,b_{k+1}\,\}-b_{k}\bigr]_{+},$ (2)
where $\cdot]_{+}\equiv\max\{\cdot,0\}$. The total tax revenue collected by the government is the summation of taxes from all agents. The government uses this tax revenue in three ways: investing in the construction of public service buildings, which are modeled as government-owned firms, distributing it to households as a universal basic income (UBI), or reserving it for the future.
@@ 4.4 Central Bank and Financial System
The central bank accepts deposits, provides loans, and implements monetary policy. In the first month of each year, it sets the policy interest rate according to a modified Taylor rule, which is a widely used monetary policy targeting rule (Galí and Gertler, 2007; Dawid and Gatti, 2018). Formally:
$\hat{r}=\max\bigl(r^{n}+\pi^{t}+\alpha(\pi-\pi^{t})+\beta(Y-Y^{n}),\,0\bigr),$ (3)
where $\hat{r}$ is the policy rate set by the central bank, $r^{n}$ is the long-run natural interest rate, $\pi^{t}$ is the target inflation rate, $\pi$ is the GDP inflation in the last period, $Y$ is the actual output (GDP), and $Y^{n}$ is the potential output as measured by a linear trend. Parameters $\alpha$ and $\beta$ capture the central bank’s responsiveness to inflation and output gap, respectively.
@@ 5 Experiments
To address this gap, we compile a checklist of canonical macroeconomic “stylized facts” from the broad economics literature (Blanchard and Fischer, 1989; Williamson, 2014; Jordà et al., 2017; Axtell and Farmer, 2025), many of which have previously been used to evaluate other ABM models. We then use this checklist to assess SimCity’s ability to reproduce established patterns and to simulate new ones under novel economic shocks. Taking advantage of SimCity’s rich framework, we aim to answer the following research questions:
@@
5.0.1 Experiment Setup
We simulate an economy with a maximum of 1000 households. Simulation proceeds in two phases. During the initial phase, which lasts 36 steps (3 years), new households gradually populate the city. During the second phase, which lasts 144 steps (12 years, 48 quarters), the population remains fixed and we observe the simulation under steady-state conditions. We use gpt-4o-mini for regular reasoning and gpt-4 as the vision language model for establishment decision-making. Both are provided by Azure OpenAI API [^1] https://portal.azure.com/ with default sampling parameter. We provide detailed examples of prompts and responses in Appendix E.3.
@@ 5.0.1 Experiment Setup
Figure 3: Beveridge Curve and other macroeconomic emergence from SimCity.
@@ 5.1 Macroeconomic Emergence (RQ1)
The Beveridge Curve captures the negative relationship between the job vacancy rate, which is the number of unfilled job openings as a proportion of total job postings, and the unemployment rate (Blanchard and Diamond, 1989; Lengnick, 2013). SimCity successfully exhibits this relationship, whereas other frameworks in comparison lack the firm module needed to explicitly model job vacancies.
Figure 4: Estimated Price Elasticity of Demand (PED) across different product categories in the SimCity environment, obeying the Law of Demand and commonsense notions of necessity.
@@ 5.2 Simulation Robustness (RQ2)
An important question for LLM-driven simulations is whether the observed economic phenomena remain stable and reproducible across runs. To address this, we run three sets of experiments with the same hyper-parameters but different random seeds, and the results demonstrate that the observed regularities are robust. Experiment results can be found at Figure 7 in Appendix B.1.
Figure 5: GDP & population curves, and map changes during the move-in phase.
@@ 5.3 Urban Expansion During Move-In (RQ3)
Figure 6: External impulse (at year 15) does not significantly affect tendency of prices of goods.
5.4 Case Study: Exogenous Price Shock (RQ4)
To investigate whether our framework can be used to analyze the effects of sudden external shocks to the economy, we separately apply a price-down and a price-up impact on the city. At the beginning of year 15, we randomly select 7 of 44 goods in two different simulations and apply a 50% price-down and a price-up impulse separately. We then let the simulation run for 6 years. Conceptually, one can think of these sudden, ephemeral changes in price levels of the selected goods as coming from exogenous shocks in the world market. Take electricity prices as an example, which is one of the shock goods in our simulations, it could come from a significant but brief conflict among energy-producing countries.
Price-down impulse response is shown as Figure 6 and price-up at Figure 8 (Appendix B.2). The experiment shows that the good’s price level gradually returns to equilibrium after an exogenous shock. Due to price stickiness, the full effects of the exogenous shock take some time to develop. Importantly, note that dotted lines in the figure represent long-run average prices of goods. Since the shock is one-off and the fundamentals of the economy did not change, prices revert close to their original levels in the long-run, despite temporal dislocation. This stylized experiment showcases SimCity’s ability to produce qualitative predictions in novel economic scenarios.

### v3

@@ Abstract
Large Language Models (LLMs) open new possibilities for constructing realistic and interpretable macroeconomic simulations. We present SimCity, a multi-agent framework that leverages LLMs to model an interpretable macroeconomic system with heterogeneous agents and rich interactions. Unlike classical equilibrium models that limit heterogeneity for tractability, or traditional agent-based models (ABMs) that rely on hand-crafted decision rules, SimCity enables flexible, adaptive behavior with transparent natural-language reasoning. Within SimCity, four core agent types (households, firms, a central bank, and a government) deliberate and participate in a frictional labor market, a heterogeneous goods market, and a financial market. Furthermore, a Vision–Language Model (VLM) determines the geographic placement of new firms and renders a virtual city map, allowing us to study both macroeconomic regularities and urban expansion dynamics within a unified environment. To evaluate the framework, we compile a checklist of canonical macroeconomic phenomena, including price elasticity of demand, Engel’s Law, Okun’s Law, the Phillips Curve, etc., and show that SimCity effectively and robustly exhibits these empirical patterns. Using our framework, we study novel economic shocks that are difficult to analyze in existing models and examine their macroeconomic implications, yielding economically coherent results.
@@ 1 Introduction
For the past two decades, dynamic stochastic general equilibrium (DSGE) models have been the predominant paradigm for studying aggregate economic behavior (Blanchard, 2009; Glandon et al., 2023). While mathematically elegant, DSGE relies on the assumption of simplified representative agents solving explicit optimization problems, which restricts heterogeneity and richer behavioral dynamics (Sergi, 2018; Vines and Wills, 2020; Storm, 2021). Agent-based models (ABMs) offer a bottom-up alternative that accommodates heterogeneity but typically rely on hand-crafted decision or learning rules (Dilaver et al., 2018). Advancing the ABM tradition with recent developments in LLMs, we propose SimCity, a multi-agent macroeconomic simulation framework that enables flexible, adaptive behavior with transparent natural language reasoning, and a virtual city environment.
SimCity models the economy as the interaction of four types of agents: households, firms, a central bank, and a government. To incorporate realistic heterogeneity, households and firms are instantiated as families of agents with rich variation in preferences, abilities, and other background, whereas the central bank and government are modeled as single institutional agents (Blanchard, 2025). Agents are implemented through an LLM-based module that integrates environmental observations, traits, and structured memory to generate reasoning, planning, and decision-making. Following the agent–interaction–environment paradigm (Wooldridge, 2009), we construct a simulation environment that features a frictional labor market, heterogeneous goods markets, and core financial interactions. The environment is augmented by a visualized map (Figure 1) to provide spatial context.
Figure 1: The SimCity framework. Left: A visualized map illustrating the urban layout across three building categories. Households reside in living buildings and engage in labor within productive buildings. City evolution is depicted in Figure 5. Right: Diagram of multi-agent interactions. Gray arrows denote the flow of capital and goods. Mechanisms of interactions are detailed in Section 4.
@@
1. LLM-driven macroeconomic simulation with rich interactions. We integrate LLMs into economic simulation by modeling four distinct economic roles, households, firms, a central bank, and a government, as LLM-driven agents. This enables agent heterogeneity and rich economic interactions.
@@
3. Systematic evaluation. We compile a checklist of canonical macroeconomic phenomena and regularities and use it to demonstrate that our framework exhibits established macroeconomic patterns. We further study responses to novel economic shocks, including exogenous price changes and shifts in production structure, highlighting our framework’s flexibility in analyzing scenarios that are difficult to capture in existing models.
@@ 2 Related Work
Simulator AI E. LEN CATS EconA. SimCity
@@
Tax Type Labor $\times$ $\times$ Labor Labor
&VAT
@@
Table 1: Comparison of features between frameworks listed in Section 5.
DSGE and ABM offer contrasting paradigms for macroeconomic modeling: DSGE emphasizes stylized optimizing agents and market-clearing equilibrium (Stokey et al., 1989; Ljungqvist and Sargent, 2018), while ABM models heterogeneous agents with rule-based behaviors (Tesfatsion and Judd, 2006). While DSGE remains dominant (Woodford, 2009; Negro et al., 2013), recent crises have renewed interest in ABMs (Stiglitz, 2018; Borsos et al., 2024). Our framework combines key advantages of both approaches by enabling heterogeneous agents with flexible, LLM-based reasoning while preserving structured economic interactions.
Li et al. (2024) introduces EconAgent as a recent attempt at LLM-driven macroeconomic simulation. Its framework restricts agent-environment interactions to two simplified variables: consumption share and work propensity, which constrains the range of economic activities that can be simulated. Our work extends this approach by incorporating other key economic elements, including firms as LLM-driven agents, heterogeneous goods markets, enriched financial activities, and a taxation system. Thus, our framework allows the exploration of a wider range of macroeconomic phenomena. Another contemporary work Mi et al. (2025) is proposed as a testbed for agent performance evaluation while we focus on urban development and macroeconomic laws.
@@ 3.1 Environment
In phase 2 (the development phase), the population is fixed while urban developments such as firm creation, investment, and spatial reallocation continue to occur. We design this phase to study macroeconomic regularities, as the system operates in a stable state that is less susceptible to the transient noise present in the expansion period.
Phase 1 lasts 36 steps, and phase 2 lasts 144 steps in a typical simulation. Each simulation step represents one month. The order of agents’ actions within each step is detailed in Appendix D.2. Agents can only access information from previous steps.
@@
We develop a web-based render module to visualize the urban expansion as shown in Figure 1. All buildings are displayed on the rendered map. The geographic placement of new firms is decided by a Visual-Language Model (VLM). For technical details of this module, please refer to Appendix D.3.
@@ 3.2 Interaction Protocol
We leverage the common-sense reasoning capabilities of LLMs to act as human-like, heterogeneous agents. Details about the prompts used are provided in Appendix E.1. Agents interact with the environment by means of function calling. The framework loads all the operations that the agent can execute and appends formatted function names along with their descriptions to the prompt. The LLM returns the actions to be taken and their parameters in JSON format. The framework will execute after a verification. For detailed examples, see Appendix E.3.
@@ 4.1 Households
Each month, households receive two types of information: a personal report summarizing income, expenditures, and other status changes during the current period, and a citywide update containing goods prices, labor market conditions, housing availability, returns on investment, and interest rates. Additionally, households may receive offers for vacancies that align with their skills.
@@ 4.2 Investment Pool and Firms
Each firm in SimCity produces a single type of good. It hires households as workers, invests in capital to improve productivity, and transforms input goods into its specialized output, which is then sold in the market. Firms are instantiated from templates that specify skill requirements of job positions in the firm and the input-output relations governing its production process, where a firm may take possibly more than one input good for its production. The effectiveness of resource utilization (i.e., the ratio at which input goods are transformed into the output good) is determined by a firm-specific Cobb-Douglas production function (Cobb and Douglas, 1928):
@@
where $A$ is total factor productivity, $K_{i}$ is the i-th firm’s capital stock, $L_{i}$ is the effective labor supply, defined as the amount of labor input adjusted for the match between job skill requirements and employees’ actual skills.
@@ 4.3 Government
$t_{i}=\sum_{k=1}^{K}r_{k}\,\max\bigl(\,\min(\,z_{i},\,b_{k+1}\,)-b_{k},0\,\bigr).$ (2)
The total tax revenue collected by the government is the sum of taxes from all agents. The government uses this tax revenue in three ways: investing in the construction of public service buildings, which are modeled as government-owned firms, distributing it to households as a universal basic income (UBI), or reserving it for the future.
@@ 4.4 Central Bank and Financial System
The central bank accepts deposits, provides loans, and implements monetary policy. In the first month of each year, it sets the policy interest rate according to a modified Taylor rule, which is widely used in monetary policies. (Galí and Gertler, 2007; Dawid and Gatti, 2018). Formally:
$\hat{r}=\max\bigl(r_{\text{natural}}+\pi_{\text{target}}+\alpha(\pi-\pi_{\text{target}})+\beta(Y-Y_{\text{natural}}),\,0\bigr),$ (3)
where $\hat{r}$ is the policy rate set by the central bank, $r_{\text{natural}}$ is the long-run natural interest rate, $\pi_{\text{target}}$ is the target inflation rate, $\pi$ is the GDP inflation in the last period, $Y$ is the actual output (GDP), and $Y_{\text{natural}}$ is the potential output as measured by a linear trend. Parameters $\alpha$ and $\beta$ capture the central bank’s responsiveness to inflation and output gap, respectively.
@@ 5 Experiments
Regularity AI-E. LEN CATS EconA. SimCity
Phillips Curve / $\times$ $\times$
Okun’s Law /
Beveridge Curve / $\times$ /
Price Elasticity / / / /
Engel’s Law / / / /
Invest. Volat. / / / /
Legend: ✓: Verified; $\times$: Inconsistent; /: Not available.
Table 2: Macroeconomic regularities exhibited by SimCity, details are in Section 5.
footnotetext: The paper by Lengnick (2013) claims to have verified the Beveridge Curve, but we are unable to reproduce it with https://github.com/newwayland/baseline-economy
@@
To address this gap, we compile a checklist of canonical macroeconomic “stylized facts” from the broad economics literature (Blanchard and Fischer, 1989; Williamson, 2014; Jordà et al., 2017; Axtell and Farmer, 2025), many of which have previously been used to evaluate other ABM models. We then use this checklist to assess SimCity’s ability to exhibit established patterns and to simulate new ones under novel economic shocks. Taking advantage of SimCity’s rich framework, we aim to answer the following research questions:
@@
5.1 Experiment Setup
We simulate an economy with a maximum of 1000 households. Each simulation proceeds in two phases. During the initial phase, which lasts 36 steps (3 years), new households gradually populate the city. During the second phase, which lasts 144 steps (12 years), the population remains fixed, and we observe the simulation under steady-state conditions. We use gpt-4o-mini for regular reasoning and gpt-4 as the vision language model for establishment decision-making. Both are provided by Azure OpenAI API [^1] https://portal.azure.com/ with default sampling parameters. We provide detailed examples of prompts and responses in Appendix E.3. We also explore the robustness of SimCity across alternative LLMs in Appendix B.1.
@@ 5.2 Macroeconomic Emergence (RQ1)
Figure 3: Macroeconomic emergences from SimCity: Beveridge Curve, Engels Law, and business cycle volatility. Business cycle volatility refers to the well-established phenomenon that investment volatility is significantly higher than consumer spending volatility across economic fluctuation (Smets and Wouters, 2007).
@@
The Beveridge Curve captures the negative relationship between the job vacancy rate, which is the number of unfilled job openings as a proportion of total job postings, and the unemployment rate (Blanchard and Diamond, 1989). SimCity successfully exhibits this relationship, whereas the previous LLM-driven ABM, EconAgent, lacks the firm module needed to model job vacancies.
@@
Figure 4: Estimated Price Elasticity of Demand (PED) across different product categories in the SimCity environment, obeying the Law of Demand and commonsense notions of necessity. $E=\frac{\Delta\log Q}{\Delta\log P}$ is the elasticity defined in Section 5.2.
@@ 5.3 Simulation Robustness (RQ2)
An important question for LLM-driven simulations is whether the observed economic phenomena remain stable and reproducible across runs. To address this, we run three sets of experiments with the same hyperparameters but different random seeds, and the results demonstrate that the observed regularities are robust. Experiment results can be found at Figure 8 in Appendix B.1.
Figure 5: GDP & population curves, and map changes during the move-in phase. Another example is given in Appendix B.4.
@@ 5.4 Urban Expansion During Move-In (RQ3)
5.5 Responses to Exogenous Shocks (RQ4)
Price shock
To investigate whether our framework can be used to analyze the effects of sudden external shocks to the economy, we separately apply a price-down and a price-up impact on the city. At the beginning of year 15, we randomly select 7 of 44 goods in two different simulations and apply a 50% price-down and a price-up impulse separately. We then let the simulation run for 6 years. Conceptually, one can think of these sudden, ephemeral changes in price levels of the selected goods as coming from exogenous shocks in the world market. Take electricity prices as an example, which is one of the shock goods in our simulations; it could come from a significant but brief conflict among energy-producing countries.
Figure 6: An external price impulse at year 15 shows that prices maintain their long-term tendency, demonstrating price stickiness.
Impulse responses to price-down shock are shown as Figure 6 and price-up at Figure 10 (Appendix B.2). The experiment shows that the good’s price level gradually returns to equilibrium after an exogenous shock. Due to price stickiness, the full effects of the exogenous shock take some time to develop. Importantly, note that dotted lines in the figure represent long-run average prices of goods. Since the shock is one-off and the economy’s fundamentals did not change, prices revert to their original levels in the long run, despite the temporary dislocation. This stylized experiment showcases SimCity’s ability to produce qualitative predictions in novel economic scenarios.
Figure 7: An automaton shock leads to an income gap between benefited agents and others.
Automation Shock
To further demonstrate the ability of our framework to simulate novel economic scenarios, we consider a stylized automation shock where physical labor is entirely replaced by information technology (IT) skills (e.g., drivers become obsolete as cars become fully autonomous). We implement this shock by altering the skill requirements in production midway through the simulation (denoted as “0 to Event” in Figure 7). As Figure 7 shows, following the shock, a substantial and widening income gap emerges between workers with strong labor skills and those with strong IT skills. This divergence persists as the economy expands, highlighting the distributional consequences of skill-biased technological change. Shaded areas represent 95% confidence intervals.